### v2

@@ 2.1 Data Bug
[…]
@@ 3.5 Determining the Symptoms
  • Z-Score analysis for weights and biases: Computing mean ($\mu$) and standard deviation ($\sigma$) identify data points with high Z-score, $|Z|>3$ (i.e., $|X-\mu|>3\sigma$tukey1977exploratory .
  • Interquartile Range (IQR) for gradients: Calculating 25th ($Q_{1}$) and 75th ($Q_{3}$) percentiles to identify values outside $[Q_{1}-1.5\times IQR,Q_{3}+1.5\times IQR]$ leys2013detecting .
  • Distribution shape analysis: Measuring skewness ($>1$) and kurtosis
@@
By capturing the above statistics from model parameters, we attempted to detect consistent statistical patterns that could serve as reliable indicators of data bugs in the DL models.
Finally, we analyzed statistics and averaged them across models and layers where these patterns appeared consistently in the buggy models. We then compared these against our bug-free models, which provided essential reference distributions of normal parameter behavior. These bug-free models were crucial for establishing reliable statistical thresholds and understanding what constitutes expected parameter distributions in properly trained models. By contrasting the buggy models’ statistics against these reference distributions, we could precisely identify which deviations represented genuine data quality issues rather than acceptable variations. These thresholds were also validated against validation datasets to confirm their effectiveness in detecting known data quality issues. This process resulted in a set of reliable statistical criteria that can help identify symptoms and their prevalence in model training.
@@ 3.7 Post-Hoc Analysis
This produces a scalar $score_{j}$ for each token $j$, representing how much attention that token receives from all other tokens on average. When calculated for all tokens, this results in a vector of importance scores $[score_{1},...,score_{n}]$, which measures each token’s relative importance in the context.
@@ 4.1.1 Impact of Data Quality
(a) Near-Zero Biases: When we trained our models using buggy data, we observed near-zero biases in several layers. Normal learning is characterized by well-defined bias values (e.g., $\geq$0.5), while near-zero biases (e.g., $<$ 0.01) might suggest learning difficulties neal2018modern . The bias parameter introduces non-linearity in the network and contributes to the learning process zhang2024bias ; williams2024expressivity . Thus, near-zero bias values might indicate a lack of non-linearity in the neural network. As shown in Table 4, 76.67% of the models trained on data with label noise and 50.00% of the models trained on data with concept drift demonstrated near-zero biases. Our manual analysis shows that $\approx$20of their layers were affected, especially the attention and output layers, which exhibited near-zero biases ($<$ 0.01) due to data bugs. In contrast, unaffected layers (e.g., embedding layer) maintained normal bias values ($\geq$ 0.5) and had regular parameter updates during backpropagation. These disparities in bias values across different layers indicate inconsistency in the learning capabilities of the layers ye2017importance . We also observed performance deterioration in the models with near-zero biases. As shown in Table 5, LineVul witnessed an 8.12%-8.39% decline across various performance metrics. CodeBERT exhibited similar behaviours for data bugs, with metrics decreasing by $\approx$6%.
@@ 4.1.2 Impact of Missing Preprocessing Operations
(b) Extreme Bias Values: When we trained the models without proper preprocessing, we observed a high range in their bias values. The absence of stop word removal had the highest impact, affecting 56.67% of LineVul and 63.33% of CodeBERT models. Our analysis revealed that $\approx$60of their layers were affected, showing widely dispersed bias values ranging between 0.03 and 1.97, significantly different from those of error-free models: (e.g., 0.34 to 0.72). Contained bias values indicate stable parameter distribution and learning patterns during model training williams2024expressivity . We also found that models with extreme bias values suffer from a substantial performance degradation, causing a 13.77%-17.32% and 7.72%-13.12% reduction in various performance metrics for our two baseline techniques (LineVul, CodeBERT).
(c) Skewed Parameter Distributions: When we trained our models without appropriate preprocessing, we also observed skewed weight distributions across the model’s layers. The absence of sequence length normalization had the highest impact, affecting 73.33% of LineVul and 83.33% of CodeBERT models. Our analysis revealed that 80% of their layers were affected, particularly embedding, attention, and dense layers. In models trained on clean code data, weights exhibited moderate skewness (0.4) and kurtosis (1.3), indicating balanced distributions. However, models trained without preprocessing showed high skewness ($>1$) and extreme kurtosis ($|k|\gg 3$), deviating significantly from clean data distributions. We also noticed a performance degradation in the affected models, causing a 10.16%-13.01% and 5.39%-10.51% reduction in various performance metrics for our two baseline techniques (LineVul, CodeBERT).
@@ 4.2.2 Impact of Missing Preprocessing Operations
(b) Skewed Bias Distributions: We also observed significant skewness in the bias distribution of our models trained without appropriate preprocessing. As shown in Table 10, 53.33% of DCCNN and 60.00% of CodeBERT+MLP models trained without stop word removal demonstrated skewed bias distributions. Our analysis of these models revealed that 50-55% of their layers were affected, particularly the convolutional layers. These layers had highly skewed bias values (e.g., $>1$) with moderate kurtosis (e.g., $|k|>3$), significantly different from the balanced distributions observed in models trained on properly preprocessed data (skewness: 0.34, kurtosis: 1.82). These skewed bias distributions also had a substantial impact on the performance of the models, causing 6.49%-7.17% and 5.75%-6.47% performance drops in various metrics (DCCNN, CodeBERT+MLP).
@@ 4.4.2 Statistical Signifiance Tests
[…]
@@
[…]
@@ 6 Related Work
Recent studies have also analyzed the impact of data quality issues on deep learning models for software engineering tasks. Wu et al. wu2021data investigated mislabeled instances in five publicly available datasets commonly used for security bug report prediction, including Chromium, Ambari, Camel, Derby, and Wicket, where they identified 749 security bug reports incorrectly labelled as non-security. They also reported significant performance improvements for a retrained model that was trained on correctly labelled data. Tantithamthavorn et al. tantithamthavorn2015impact analyzed over 3,900 issue reports from Apache projects and demonstrated that label noise can significantly impact the recall of a model. They also reported a significant improvement in the recall ($\approx$60%) for a retrained model on cleaned datasets. Kim et al. kim2011dealing evaluated the impact of intentionally injected noise into the datasets for defect prediction models through controlled experiments. They found that when $\approx$30of the data was mislabeled, the performance of defect prediction models decreased significantly, which highlights the models’ sensitivity to label noise. Fan et al. fanimpact investigated mislabeled changes in just-in-time defect prediction datasets and found that certain labelling approaches can lead to performance reduction of up to 5%. Xu et al. xu2023data leveraged adversarial learning to improve the data quality of obsolete comment datasets, which led to an improvement in the accuracy of multiple existing models by 18.1%. Croft et al. croft2023data conducted a systematic analysis of data quality in software vulnerability datasets and found out that 71% of the labels are incorrect and 17-99% of data points are duplicated across four state-of-the-art datasets. Furthermore, they also analyzed the impact of the data quality issues on vulnerability detection. They found that the model’s performance dropped by 65% when trained on clean data, which shows how the duplication and mislabelling of the data points can lead to inflated results.

### v3

@@ 2.1 Data Bug
$\exists\ \mathcal{E}\in\mathcal{P}\ \text{such that}\ \mathcal{E}(X_{\text{raw}})\neq X_{\text{expected}},$
@@
Out of Distribution Data: Out-of-distribution (OOD) data refers to data that does not come from the probability distribution of the training data. A trained model is assumed to generalize well to new data from the training distribution. OOD data violates this assumption and could lead to poor model performance.
Formally, given a training distribution $P_{\text{in}}(X,Y)$ and a test distribution $P_{\text{out}}(X,Y)$, a data point $(x,y)$ is considered out-of-distribution if it is drawn from $P_{\text{out}}$ and $P_{\text{out}}\neq P_{\text{in}}$.
The core challenge with OOD data is that a model, having learned patterns specific to the training distribution, may produce overconfident or incorrect predictions for OOD samples, as it has not been exposed to similar examples during training. While OOD characteristics can arise from a natural domain shift, they become a data quality issue when a defect or an error causes a significant deviation from the training distribution. For example, consider a model trained to classify images of cars. If the model is trained only on images of a specific type of car (e.g., sedans), a new set of images of another type of car (e.g., pickup trucks), can be considered as out-of-distribution data due to a natural distributional shift. However, if the new images contain corrupt data (e.g., solid black images) or incorrect labels, this out-of-distribution deviation represents a data quality issue caused by errors in preprocessing or manual annotation.
@@ 3.5 Determining the Symptoms
[…]
  • Interquartile Range (IQR) for gradients: Calculating 25th ($Q_{1}$) and 75th ($Q_{3}$) percentiles to identify values outside $Q_{1}-1.5\times IQR,Q_{3}+1.5\times IQR]$ leys2013detecting .
  • Distribution shape analysis: Measuring skewness (>1>1) and kurtosis
@@
These observed statistical deviations are the ”symptoms” of data quality issues. The correlation between a symptom and a specific data quality issue is established by observing consistent patterns in our controlled experiments. For instance, when we introduce a bug like class imbalance, a specific statistical pattern (e.g., high gradient skewness) consistently emerges across multiple training runs. Since this pattern is absent in our bug-free baseline models, we can directly correlate the bug with its statistical symptom.
Finally, we analyzed statistics and averaged them across models and layers where these patterns appeared consistently in the buggy models. We then compared these against our bug-free models, which provided essential reference distributions of normal parameter behavior. These bug-free models were crucial for establishing reliable statistical thresholds and understanding what constitutes expected parameter distributions in properly trained models. Therefore, the thresholds are the specific statistical boundaries derived from this comparative analysis (e.g., a Z-score >> 3 or skewness >> 1), which signal a high probability of a data quality issue when exceeded. These thresholds were also validated against validation datasets to confirm their effectiveness in detecting known data quality issues. This process resulted in a set of reliable statistical criteria that can help identify symptoms and their prevalence in model training.
@@ 3.7 Post-Hoc Analysis
This produces a scalar $score_{j}$ for each token $j$, representing how much attention that token receives from all other tokens on average. When calculated for all tokens, this results in a vector of importance scores $score_{1},...,score_{n}]$, which measures each token’s relative importance in the context.
@@ 4.1.1 Impact of Data Quality
(a) Near-Zero Biases: When we trained our models using buggy data, we observed near-zero biases in several layers. Normal learning is characterized by well-defined bias values (e.g., $\geq$0.5), while near-zero biases (e.g., << 0.01) might suggest learning difficulties neal2018modern . The bias parameter introduces non-linearity in the network and contributes to the learning process zhang2024bias ; williams2024expressivity . Thus, near-zero bias values might indicate a lack of non-linearity in the neural network. As shown in Table 4, 76.67% of the models trained on data with label noise and 50.00% of the models trained on data with concept drift demonstrated near-zero biases. Our manual analysis shows that $\approx$20% of their layers were affected, especially the attention and output layers, which exhibited near-zero biases (<< 0.01) due to data bugs. In contrast, unaffected layers (e.g., embedding layer) maintained normal bias values ($\geq$ 0.5) and had regular parameter updates during backpropagation. These disparities in bias values across different layers indicate inconsistency in the learning capabilities of the layers ye2017importance . We also observed performance deterioration in the models with near-zero biases. As shown in Table 5, LineVul witnessed an 8.12%-8.39% decline across various performance metrics. CodeBERT exhibited similar behaviours for data bugs, with metrics decreasing by $\approx$6%.
@@ 4.1.2 Impact of Missing Preprocessing Operations
(b) Extreme Bias Values: When we trained the models without proper preprocessing, we observed a high range in their bias values. The absence of stop word removal had the highest impact, affecting 56.67% of LineVul and 63.33% of CodeBERT models. Our analysis revealed that $\approx$60% of their layers were affected, showing widely dispersed bias values ranging between 0.03 and 1.97, significantly different from those of error-free models: (e.g., 0.34 to 0.72). Contained bias values indicate stable parameter distribution and learning patterns during model training williams2024expressivity . We also found that models with extreme bias values suffer from a substantial performance degradation, causing a 13.77%-17.32% and 7.72%-13.12% reduction in various performance metrics for our two baseline techniques (LineVul, CodeBERT).
(c) Skewed Parameter Distributions: When we trained our models without appropriate preprocessing, we also observed skewed weight distributions across the model’s layers. The absence of sequence length normalization had the highest impact, affecting 73.33% of LineVul and 83.33% of CodeBERT models. Our analysis revealed that 80% of their layers were affected, particularly embedding, attention, and dense layers. In models trained on clean code data, weights exhibited moderate skewness (0.4) and kurtosis (1.3), indicating balanced distributions. However, models trained without preprocessing showed high skewness (>1>1) and extreme kurtosis ($|k|\gg 3$), deviating significantly from clean data distributions. We also noticed a performance degradation in the affected models, causing a 10.16%-13.01% and 5.39%-10.51% reduction in various performance metrics for our two baseline techniques (LineVul, CodeBERT).
@@ 4.2.2 Impact of Missing Preprocessing Operations
(b) Skewed Bias Distributions: We also observed significant skewness in the bias distribution of our models trained without appropriate preprocessing. As shown in Table 10, 53.33% of DCCNN and 60.00% of CodeBERT+MLP models trained without stop word removal demonstrated skewed bias distributions. Our analysis of these models revealed that 50-55% of their layers were affected, particularly the convolutional layers. These layers had highly skewed bias values (e.g., >1>1) with moderate kurtosis (e.g., $|k|>3$), significantly different from the balanced distributions observed in models trained on properly preprocessed data (skewness: 0.34, kurtosis: 1.82). These skewed bias distributions also had a substantial impact on the performance of the models, causing 6.49%-7.17% and 5.75%-6.47% performance drops in various metrics (DCCNN, CodeBERT+MLP).
@@ 4.4.2 Statistical Signifiance Tests
$H_{0}$ $:\text{The presence of data quality issues and model}$
@@
$H_{1}$ $:\text{The presence of data quality issues and model}$
@@ 5 Discussion
5.1 Impact of Data Quality Issues for Generation Tasks
Generation tasks have gained significant attention in the past few years, thanks to the emergence of large language models. As shown in existing studies yang2022dlinse , these tasks are the second most prevalent type of task in software engineering. To understand the impact of data bugs on a generation task, we conduct a prototype experiment that investigates how data quality and preprocessing issues affect code summarization. We utilized two well-known datasets, FunCom and TLCodeSum, which have been used in prior work mu2023developer ; gao2023makes ; gros2020code and are known to contain data quality issues shi2022we ; vitale2025optimizingdatasetscodesummarization .
To address these issues from the datasets, we employed two distinct techniques from prior work shi2022we ; vitale2025optimizingdatasetscodesummarization . First, we used SIDE-based filtering, which measures the coherence between code snippets and comments using contrastive learning, and filters out instances with low SIDE scores. Additionally, we employed CAT, a rule-based preprocessing tool that leverages an existing taxonomy shi2022we to identify the comment-related and code-related noise in the dataset. The tool then applies two primary strategies to clean the data. The first strategy is to fix minor issues in comments by removing inappropriate or noisy content, such as boilerplate text and block-level comments. The second strategy is to completely drop instances where the code itself is structurally problematic, for example, removing code-comment pairs for simple getter and setter methods. We use CodeT5+ as the base model, following the approach employed in existing studies vitale2025optimizingdatasetscodesummarization .
We trained three variants of CodeT5+ to evaluate the effects of data quality and preprocessing. These variants are described below:
  • Clean: The model was trained on datasets that had been carefully cleaned using both the SIDE filtering tool and the CAT preprocessing tool.
  • Buggy Data: The model was trained on the datasets without SIDE filtering. This allowed us to assess the impact of raw, unfiltered data containing inherent quality issues (e.g., outdated code-comment pairs).
  • Missing Preprocessing: The model was trained on datasets that were not preprocessed using the CAT preprocessing tool.
When we trained the models with buggy data, we observed a significant difference in their gradient behaviours. Gradient histograms of the models trained on clean data demonstrated tightly concentrated values. For example, a representative gradient distribution from the clean dataset shows values predominantly in the range of approximately $1.12\times 10^{-5}$ to $7.44\times 10^{-5}$. This central clustering indicates stable and consistent parameter updates during backpropagation. In contrast, models trained on buggy data exhibit significant instability in their gradients. Their gradient distributions are often wider and irregular, with a noticeable presence of outliers. For instance, a gradient histogram from the low-quality dataset shows a broader distribution (e.g., 0.000157-0.000157 to $8.54\times 10^{-6}$). This dispersed distribution indicates less stable gradient updates and potential issues during the optimization process. We observed this behaviour in classification-based tasks as well, suggesting the generalizability of our findings.
The absence of a preprocessing step also significantly impacts a model’s training process. Our analysis of parameter distributions reveals the effect of missing preprocessing on model training. Models trained with adequate preprocessing demonstrate well-defined and stable weight distributions. A representative parameter histogram from our clean model shows that values are evenly distributed within a contained range (e.g., from 0.098 to 0.383). This indicates that the model has learned a structured and balanced set of weights, which is essential for effective generalization. Conversely, a model trained without adequate preprocessing shows a broader and more dispersed weight distribution, with values spanning a wider range (e.g., from 0.061 to 0.766). This pattern, also seen in classification tasks, further supports the generalizability of these findings.
@@ 6 Related Work
Recent studies have also analyzed the impact of data quality issues on deep learning models for software engineering tasks. Wu et al. wu2021data investigated mislabeled instances in five publicly available datasets commonly used for security bug report prediction, including Chromium, Ambari, Camel, Derby, and Wicket, where they identified 749 security bug reports incorrectly labelled as non-security. They also reported significant performance improvements for a retrained model that was trained on correctly labelled data. Tantithamthavorn et al. tantithamthavorn2015impact analyzed over 3,900 issue reports from Apache projects and demonstrated that label noise can significantly impact the recall of a model. They also reported a significant improvement in the recall ($\approx$60%) for a retrained model on cleaned datasets. Kim et al. kim2011dealing evaluated the impact of intentionally injected noise into the datasets for defect prediction models through controlled experiments. They found that when $\approx$30% of the data was mislabeled, the performance of defect prediction models decreased significantly, which highlights the models’ sensitivity to label noise. Fan et al. fanimpact investigated mislabeled changes in just-in-time defect prediction datasets and found that certain labelling approaches can lead to performance reduction of up to 5%. Xu et al. xu2023data leveraged adversarial learning to improve the data quality of obsolete comment datasets, which led to an improvement in the accuracy of multiple existing models by 18.1%. Croft et al. croft2023data conducted a systematic analysis of data quality in software vulnerability datasets and found out that 71% of the labels are incorrect and 17-99% of data points are duplicated across four state-of-the-art datasets. Furthermore, they also analyzed the impact of the data quality issues on vulnerability detection. They found that the model’s performance dropped by 65% when trained on clean data, which shows how the duplication and mislabelling of the data points can lead to inflated results.